PLEASE HELP TO CALCULATE THE REMAINING 4 VALUES. SHOW WORK

Determining the Dimensions of a Molecule

Volume of PURE oleic acid   (show calculations below)                       0.027ml___ mL

0.027ml = 12 drops * 0.0152ml (one drop)

Number of oleic acid molecules in the slick (show calculations)         1.62594*1022

0.027ml * (6.022*1023) = 1.62594*1022

Surface area of the oil slick (show calculations)                                   237.72____ cm2

3.14(8.701cm)2 = 237.72

Cross sectional area of an oleic acid molecule (show calculations)        __________ cm2

Depth of the oil slick (show calculations)                                               __________ cm

Diameter of an oleic acid molecule (show calculations)                         __________ cm

Length of an oleic acid molecule                                                             __________ cm
Answer #1

1- volume of the oleic acid molecule= volume of the pure oleic acid / total molecule

=0.027/1.625x1022=1.66x10-24mL=4/3r3

r=7.34x10-9cm

cross sectional area= r2=3.14x(7.34x10-9)2=1.69x10-16cm2

2- depth of the oil slick=volume of the oil slick/ cross sectional area

=0.027/237.72=1.14x10-4cm

3- diameter of an oleic acid molecule= 2r =2x7.34x10-9=1.468x10-8cm.

4- length of the oleic acid molecule=diameter=1.468x10-8cm